OpenStudy (anonymous):

So I'm looking at the Key Concept "the nth Root" on p. 361. I understand that When n is even that you have two real nth roots of (+)b. And when n is odd you only have one real nth root of (+)b. So I get the odd and even part about n. What I don't understand is why it only tells you there are no real nth roots of (-)b, when n is even. So my question would be; what happens when n is odd and b is negative? Does the fact that when b is negative, regardless the charge of n, mean that there will not be any real nth roots?

OpenStudy (anonymous):

good question :)

OpenStudy (anonymous):

What I don't understand is why it only tells you there are no real nth roots of (-)b, when n is even. lets first think about this

OpenStudy (anonymous):

3x3 = 9 so, \(\large \sqrt[2]{9} = 3 \)

OpenStudy (anonymous):

can u find any number which gives a negative number, when multiplied by itself ?

OpenStudy (anonymous):

I've been doing research on this and I think that it doesn't matter.

OpenStudy (anonymous):

Any number in this? Or in general?

OpenStudy (anonymous):

in general, just find some real number \(n\) wid below property :- \(\large n \times n\) = negative number

OpenStudy (anonymous):

a negative number times itself (which is negative) will be negative.. So would it be that you can have a negative -b and have roots when n is odd?? That's what im finding..

OpenStudy (anonymous):

multiplying a number by itself, fails to produce a negative numbebr. so for negative numbers, \(n\)th roots are not defined when \(n\) is 2, 4, 6...

OpenStudy (anonymous):

So what you're saying is that you can have a -b ONLY when n is odd.. RIGHT?! xD

OpenStudy (anonymous):

im saying, when \(n\) is even, and \(b\) is negative, then \(n\) th root is not defined in reals below thing doesnt exist in real numbers :- \(\large \sqrt[n]{neg}\)

OpenStudy (anonymous):

I get that. When n is odd and b is negative. You CAN have one real root. That's what my question was in the first place lol

OpenStudy (anonymous):

read ur question again, ur first question was on why there are no even nth roots for negative numbers :o

OpenStudy (anonymous):

Right cuz the book doesn't say anything about it not being possible to have roots when you have an odd nth power for a negative number. Its all good I figured it out.